A 60 kgm body has a potential energy of 135 Kcal with respect to a datum plane X and a potentiaal energy of -99 Kcal with respect to datum plane Y. If the local acceleration of gravity is standard, compute the relative position of plane X relative to plane Y. Resolve the problem if g = fps2.


1
Expert's answer
2021-01-25T04:09:34-0500

m = 60 kg

E1 = 135 Kcal = 564 840 J

E2 = -99 Kcal = -414 216 J

E = mgh

60 kg x 9.8 m/s2 x h1 = 564 840 J

h1 = 961 m

60 kg x 9.8 m/s2 x h2 = -414 216 J

h2 = -704 m

Relative position = 961 - 704 = 257